# Directional Diffusion-Style Code Editing Pre-training

### v1

@@ II-D Diffusion Models for Text Generation
Diffusion models, which predominantly operate at the continuous embedding level, have made remarkable progress in continuous domains such as image generation [13]. These models involve two key steps: forward noising process and reverse denoising process [25]. Specifically, the forward noise addition process gradually transforms the original data into pure Gaussian noise using a Markov chain, with each noise addition step represented as:
@@
Existing works applying diffusion models to discrete text data introduce an additional step in the noise addition process, requiring the conversion of discrete text W into continuous embeddings (i.e., $q_{\phi}(\mathbf{X_{0}}|\mathbf{W})=\mathcal{N}(EMB(\mathbf{W}),\alpha_{0}\mathbf{I})$). Correspondingly, the denoising process includes a trainable rounding step to convert embeddings back into discrete text (i.e., $p_{\theta}(\mathbf{W}|\mathbf{X_{0}})=\prod_{i=1}^{N}p_{\theta}(W_{i}|X_{i})$[15, 14, 26]. For example, in the upper part of Figure 2, traditional diffusion performs a left-to-right denoising process. The model progressively denoises a continuous embedding and ultimately converts the continuous embedding into discrete token representations through a rounding layer, generating the final text (e.g., ‘ABC’). Similarly, during the inference process, traditional diffusion also progressively denoises a continuous embedding. Notably, this embedding represents the entire text, and each denoising step generates a new embedding, which is then converted into all tokens of the text (i.e., in a non-autoregressive manner). As a result, every token in the text has the potential to change at each step of the inference process (e.g., from pure noise to ‘fdc’, from ‘fdc’ to ‘aGh’, etc.). This implies that achieving coherence and logical consistency across all tokens often requires many iterative steps.
@@
Our approach diverges from existing methods that apply diffusion to discrete text in two key ways. First, rather than adding noise at the continued embedding level and iteratively removing it, we simulate real-world code editing by leveraging actual code evolution paths. For example, in Figure 2, we treat the partially modified text during the evolution process as “noise” (e.g., CbDaEa), while the target of the evolution is considered the “noise-free” text (e.g., CbDbEb). This approach assigns a concrete meaning to each instance of “noise”, effectively simulating real-world editing scenarios. Second, we implement a diffusion-style process by utilizing the intermediate state data from the editing process to generate the final noise-free text, where each “denoising” step is treated as a transition from a noisy state to a noise-free state. This ensures that we can generate directly usable code in a single step—without multiple denoising iterations—at a computational cost comparable to standard basic models (e.g., in an autoregressive manner like Transformer [27], CodeT5). Specifically, the inference time complexity of traditional diffusion applied to text is about $O(TL^{2}H)$, while combining the diffusion concept with traditional text generation models results in a complexity of $O(L^{3}H)$, where $T$ is the number of diffusion steps, $L$ is the sequence length, and $H$ is the hidden layer dimension. Typically, $T$ is significantly larger than $L$ (e.g., a sequence of 200 tokens may require up to 2000 denoising steps). Thus, traditional text generation requires less inference time [14] (The detailed information in SectionVI).
@@
Figure 3: An denoising example. The process of modifying the old code (center) to fulfill the requirements outlined in the comments (top left) often necessitates changes across multiple dispersed locations. Achieving the desired evolution (right) requires incremental editing at all relevant points.
@@ III-A Model Architecture
We employ an encoder-decoder framework with an auto-regressive manner based on the Transformer [27] model for pre-training. Following the existing code editing pre-trained models [11, 8, 7], we continue pre-training DivoT5 on top of the weights of CodeT5 [9]. This enables the model to leverage CodeT5’s knowledge, offering significant advantages in code-related tasks.
@@ III-C2 Diffusion for the Evolutionary Direction
Considering the above challenges, DivoT5, designed for discrete data like code, employs discrete-data-level diffusion, where each noise represents a specific code change, reflecting the natural evolution of code. This strategy allows the entire approach to be completed in an auto-regressive manner as well as to leverage the advantages of existing code-related approaches [9, 11, 7, 8]. In the noising stage, we first introduce artificial noise into the old code to learn token-level code editing processes and enrich the context of the old code. Then, we treat the intermediate evolution process of editing as a special type of evolution noise to enhance the direction of code evolution. Noisy data (artificial or evolutionary) can be paired with the final new code to form input-output pairs, enabling the model to learn the direction of evolution from different contexts. In the denoising stage, we treat different types of noise as distinct pre-training tasks, and denoising is performed fully in an auto-regressive manner. This enables the model to consider both the contextual dependencies of code and the incremental characteristics of code evolution. Below, we discuss the noising process at the data level and the denoising process in an auto-regressive manner.
@@
Code evolution typically requires referencing an existing code, making the starting point of directional diffusion not pure Gaussian noise but an old code that needs modification ($X_{T}$). On this basis, we add artificial noise to perturb the old data ($X_{T+1}$), simulating the basic editing process and enriching the context of the old code. This approach aims to help the model grasp fine-grained code edits and accurately generate the target new code across various contexts. We use $X_{t}$ to represent an intermediate state in the evolution process, indicating that $t$ modifications are still needed to complete all editing operations. For example, in Figure 3, $X_{T-1}$ represents the state where the model has modified the block $O_{T}$ in the old code $X_{T}$ (e.g., adding the import of a package). Similarly, $X_{T-2}$ represents the state where the model has modified the block $O_{T-1}$ in $X_{T-1}$ (e.g., changing variables and implementing function bodies). To obtain the intermediate version $X_{t}$, we calculate the diff between the old code file and the new code file, and then split the diff based on the distance between them. Specifically, if transforming old code A (‘CaDaEa’) to new code B (‘CbDbEb’) involves three changes (each change representing a diff hunk that may affect several nearby lines), the intermediate code evolves incrementally: first modifying the first place to produce ‘CbDaEa‘, then applying the second change to produce ‘CbDbEa’. These intermediate versions follow the edit history sequence, progressing from the beginning of the code to its end and mirroring typical code editing practice. We treat the code that has undergone partial human modifications and still requires further modifications as a sample containing evolutionary noise. By gradually adding artificial noise and evolutionary noise (human edits) during the code evolution process, we can effectively simulate the transition of code from its evolved state (i.e., $X_{0}$) to the corrupted state (i.e., $X_{T+1}$). The process of adding noise can be formalized as follows:
@@
, where $q(X_{t}|X_{0})$ represents the process of gradually adding noise from $X_{0}$ to $X_{t}$, and $f_{i}(X_{0})$ is a $i$-th function representing the addition of noise to $X_{i-1}$ to obtain $X_{i}$. For example, in Figure 3, $f_{1}$ represents the operation of reverting the modification to the $O_{1}$ block in the evolved code, indicating that there is still a part of the code that needs to be modified. Each $X_{t}$ represents the state of the code at a specific stage of evolution, providing explicit direction that helps the model approach the final state at different modification stages (i.e., constructing the input-output pairs like $<X_{t}$, $X_{0}>$). This strategy improves the model’s adaptability to diverse contexts and enhances its ability to capture subtle changes in the code.
@@
Existing diffusion models for discrete text generation typically add and remove noise at the continuous embedding level step by step. While this can increase the diversity of the generated text, it requires more iterative steps, making the process slower compared to the mainstream auto-regressive text generation approaches, resulting in lower efficiency. For example, in the upper part of Figure 2, each denoising step in traditional diffusion models modifies the overall embedding, allowing every token to potentially change in the next step, which facilitates generating diverse text. In contrast, in autoregressive generation, once a token is generated, its content and position are fixed, and the generation continues sequentially from that point onward, facilitating semantic coherence in the output [27, 9]. Therefore, we conduct the denoising process within an auto-regressive manner similar to CodeT5, using any intermediate state $X_{t}$ from the diffusion process as input and producing the final target $X_{0}$ as output. This approach maximizes the generative performance benefits of auto-regression. The denoising process can be formalized as:
@@
, where $p_{\theta}(X_{t-1}|X_{t})$ represents the probability distribution of denoising from $X_{t}$ to $X_{t-1}$, and $p_{\theta}(X_{t-1}^{i}|X_{t-1}^{1:i-1};X_{t})$ represents the probability of generating the $i$-th token in $X_{t-1}$, given the existing tokens in $X_{t}$ and generated tokens in $X_{t-1}$. The core characteristic of directional diffusion is reflected in each generation process of the auto-regressive model, where the generation target at any step is the final evolution goal. In other words, we transform intermediate diffusion data into input-output pairs for the auto-regressive model, such as $<X_{T+1}$, $X_{0}>$, $<X_{T}$, $X_{0}>$, $<X_{T-1}$, $X_{0}>$. Under this strategy, each sample in the generation process is directional and optimized towards the final new code, thereby enhancing the model’s accuracy and consistency in generating outputs. Notably, by converting the directional diffusion process into input-output pairs for the pre-training task, our pre-trained model has the same inference cost as auto-regressive models like CodeT5 (requiring a single inference pass), which is lower compared to the higher inference cost of traditional diffusion models (requiring multiple inference passes).
@@ III-D Directional Diffusion Pretraining Tasks
DivoT5 aims to learn real-world code editing knowledge to correctly generate new evolved code from old code through four proposed pre-training tasks for directional diffusion. To better illustrate the pre-training tasks, let us assume that the model only has the old code x ($x^{1},x^{2},\cdots,x^{M}$) as the input and the new code y ($y^{1},y^{2},\cdots,y^{N}$) as the output, where $M$ represents the number of input tokens and $N$ represents the number of output tokens. Figure 3 shows an illustration of the four tasks, where the corrupted code on the left ($X_{T+1}$) represents examples of the first three tasks, and the intermediate states $X_{t}$ ($0<t\leq T$) from the old code ($X_{T}$) in the middle to the new code ($X_{0}$) on the right illustrates the input for the last task. For clarity, we refer to the input code as the old code and the output code as the new code. The first three tasks are designed to learn the basic editing processes and contextual dependencies of the code during the diffusion process, while the final task aims to reinforce the direction of code evolution. We place an additional special token at the prefix of the input sequence to distinguish between different tasks during the training process. Below we discuss the details of these tasks.
@@
This task aims to make the model be aware of which code tokens do not need to be modified (i.e., Keep Span) during the code editing process. The design of this pre-training task takes into account the token-level similarity in code editing. Since the code before and after evolution typically follows similar patterns or structures, some content may remain unchanged after editing. For example, as shown in the $X_{T+1(a)}$ of Figure 3, $[MASK1]$ represents the keywords that define the variable, which is typically not changed both before and after the edit. Therefore, we design a new technique in our editing data by masking only the tokens that remain unchanged. This allows the model to better learn the common parts frequently shared during code editing, thereby enhancing its ability to understand and edit code. Applying this task of masking tokens to a real-world code editing dataset involves not only masking the shared tokens but also generating new code that aligns with the direction of evolution.
Following existing works [28, 8], we calculate the differences between the new code and the old code. However, unlike existing techniques [11, 7], our difference information is at the token level rather than at the line level. Then, we apply the SMLM [29] task specifically to the tokens marked as “KEEP” in the difference information. The formal representation of this task is as follows:
@@
, where $i$ represents the current generation step, $y_{i}$ represents the $i$-th token in the new code $y$. The $mask_{KSM}$ represents a type of SMLM approaches, $x_{keep}$ indicates the unchanged parts of the code before and after editing, and $x_{change}$ indicates the modified parts. Since the number of unchanged tokens in this task is smaller than the total number of tokens in the old code, we mask 30% of the unchanged tokens, aiming to achieve a similar total number of masked tokens as in other tasks.
@@
This pre-training task treats the code editing process as a text generation task with evolutionary direction. We randomly mask a set of token spans in the input and sequentially predict the new code using a sequence-to-sequence framework. Based on the characteristics of the dataset and existing works [7, 8], we set the average span count to 2.5 and the overall percentage of masked tokens to be 20%. Unlike self-supervised tasks in the field of general text generation, the goal of this pre-training task is not simply to predict masked words, but rather to generate edited code in the presence of masks (i.e., prediction of the evolutionary direction). Formally, the loss of this task can be describe as follows:
@@
Our third pre-training task combines the characteristics of denoising auto-encoding with the real-world code editing scenarios. It aims to train the model to generate the new code accurately while taking into account the noise and variations that occur in old code. Compared to the previous two pre-training tasks, this task demands a stronger ability for the model to understand code editing. However, this task well aligns with real-world development scenarios because we often cannot guarantee the quality of the old code. It reflects the practical nature of code editing where the quality and effectiveness of the existing code vary. Specifically, we apply three types of noise to the old code: insertion, deletion, and replacement [8]. As shown in the code snippet $X_{T+1(c)}$ of Figure 3, in the old code, we add a $[MASK0]$ token, delete the keyword tokens like $public$, and replace specific variable names with $[MASK1]$. These modifications introduce noise and variations to simulate real-world code editing scenarios. During the generation process, the model not only needs to consider the tokens replaced by $[MASK]$, but also needs to take into account the scenario where existing tokens are deleted and the case of useless masks. Unlike existing techniques[8, 10] that utilize DAE with controlled noises to restore the old code, our approach requires the model to generate the actual new code after removing the introduced noise. The loss function for this pre-training task is as follows:

### v2

@@ II-D Diffusion Models for Text Generation
Diffusion models, which predominantly operate at the continuous embedding level, have made remarkable progress in continuous domains such as image generation [diffapp_ho2020denoising]. These models involve two key steps: forward noising process and reverse denoising process [diff_sohl2015deep]. Specifically, the forward noise addition process gradually transforms the original data into pure Gaussian noise using a Markov chain, with each noise addition step represented as:
@@
Existing works applying diffusion models to discrete text data introduce an additional step in the noise addition process, requiring the conversion of discrete text W into continuous embeddings (i.e., $q_{\phi}(\mathbf{X_{0}}|\mathbf{W})=\mathcal{N}(EMB(\mathbf{W}),\alpha_{0}\mathbf{I})$). Correspondingly, the denoising process includes a trainable rounding step to convert embeddings back into discrete text (i.e., $p_{\theta}(\mathbf{W}|\mathbf{X_{0}})=\prod_{i=1}^{N}p_{\theta}(W_{i}|X_{i})$[difflm_DiffLM_li2022diffusion, difflm_ARDiff_wu2024ar, difflm_Genie_lin2023text]. For example, in the upper part of Figure 2, traditional diffusion performs a left-to-right denoising process. The model progressively denoises a continuous embedding and ultimately converts the continuous embedding into discrete token representations through a rounding layer, generating the final text (e.g., ‘ABC’). Similarly, during the inference process, traditional diffusion also progressively denoises a continuous embedding. Notably, this embedding represents the entire text, and each denoising step generates a new embedding, which is then converted into all tokens of the text (i.e., in a non-autoregressive manner). As a result, every token in the text has the potential to change at each step of the inference process (e.g., from pure noise to ‘fdc’, from ‘fdc’ to ‘aGh’, etc.). This implies that achieving coherence and logical consistency across all tokens often requires many iterative steps.
@@
Our approach diverges from existing methods that apply diffusion to discrete text in two key ways. First, rather than adding noise at the continued embedding level and iteratively removing it, we simulate real-world code editing by leveraging actual code evolution paths. For example, in Figure 2, we treat the partially modified text during the evolution process as “noise” (e.g., CbDaEa), while the target of the evolution is considered the “noise-free” text (e.g., CbDbEb). This approach assigns a concrete meaning to each instance of “noise”, effectively simulating real-world editing scenarios. Second, we implement a diffusion-style process by utilizing the intermediate state data from the editing process to generate the final noise-free text, where each “denoising” step is treated as a transition from a noisy state to a noise-free state. This ensures that we can generate directly usable code in a single step—without multiple denoising iterations—at a computational cost comparable to standard basic models (e.g., in an autoregressive manner like Transformer [transformer2017], CodeT5). Specifically, the inference time complexity of traditional diffusion applied to text is about $O(TL^{2}H)$, while combining the diffusion concept with traditional text generation models results in a complexity of $O(L^{3}H)$, where $T$ is the number of diffusion steps, $L$ is the sequence length, and $H$ is the hidden layer dimension. Typically, $T$ is significantly larger than $L$ (e.g., a sequence of 200 tokens may require up to 2000 denoising steps). Thus, traditional text generation requires less inference time [difflm_ARDiff_wu2024ar] (The detailed information in SectionVI).
@@
Figure 3: The process of modifying the old code (center) to fulfill the requirements outlined in the comments (top left) often necessitates changes across multiple dispersed locations. Achieving the desired evolution (right) requires incremental editing at all relevant points. The blue italics aim to highlight the parts that persist throughout the entire code editing process.
@@ III-A Model Architecture
We employ an encoder-decoder framework with an auto-regressive manner based on the Transformer [transformer2017] model for pre-training. Following the existing code editing pre-trained models [lin2023cct5, coditt52022, codereviewer, liang2024condor], we continue pre-training DivoT5 on top of the weights of CodeT5 [codet52021]. This enables the model to leverage CodeT5’s knowledge, offering significant advantages in code-related tasks [coditt52022, chen2024improving, zhao2024spotting].
@@ III-C2 Diffusion for the Evolutionary Direction
Considering the above challenges, DivoT5, designed for discrete data like code, employs discrete-data-level diffusion, where each noise represents a specific code change, reflecting the natural evolution of code. This strategy allows the entire approach to be completed in an auto-regressive manner as well as to leverage the advantages of existing code-related approaches [codet52021, lin2023cct5, codereviewer, coditt52022, liang2021lyra, wang2024advanced]. In the noising stage, we first introduce artificial noise into the old code to learn token-level code editing processes and enrich the context of the old code. Then, we treat the intermediate evolution process of editing as a special type of evolution noise to enhance the direction of code evolution. Noisy data (artificial or evolutionary) can be paired with the final new code to form input-output pairs, enabling the model to learn the direction of evolution from different contexts. In the denoising stage, we treat different types of noise as distinct pre-training tasks, and denoising is performed fully in an auto-regressive manner. This enables the model to consider both the contextual dependencies of code and the incremental characteristics of code evolution. Below, we discuss the noising process at the data level and the denoising process in an auto-regressive manner.
@@
Code evolution typically requires referencing an existing code, making the starting point of directional diffusion not pure Gaussian noise but an old code that needs modification ($X_{T}$). On this basis, we add artificial noise to perturb the old data ($X_{T+1}$), simulating the basic editing process and enriching the context of the old code. This approach aims to help the model grasp fine-grained code edits and accurately generate the target new code across various contexts. We use $X_{t}$ to represent an intermediate state in the evolution process, indicating that $t$ modifications are still needed to complete all editing operations. For example, in Figure 3, $X_{T-1}$ represents the state where the model has modified the block $O_{T}$ in the old code $X_{T}$ (e.g., adding the import of a package). Similarly, $X_{T-2}$ represents the state where the model has modified the block $O_{T-1}$ in $X_{T-1}$ (e.g., changing variables and implementing function bodies). To obtain the intermediate version $X_{t}$, we compute the diff between the old and new code files using Python’s difflib library, and then split the diff into hunks based on the relative positions of the changes within the code. Specifically, if transforming old code A (‘CaDaEa’) to new code B (‘CbDbEb’) involves three changes (each change representing a diff hunk that may affect several nearby lines), the intermediate code evolves incrementally: first modifying the first position to produce ‘CbDaEa’, then applying the second change to produce ‘CbDbEa’. These intermediate versions follow the edit history sequence, progressing from the beginning of the code to its end and mirroring typical code editing practice. We treat the code that has undergone partial human modifications and still requires further modifications as a sample containing evolutionary noise. By gradually adding artificial noise and evolutionary noise (human edits) during the code evolution process, we can effectively simulate the transition of code from its evolved state (i.e., $X_{0}$) to the corrupted state (i.e., $X_{T+1}$). The process of adding noise can be formalized as follows:
@@
, where $q(X_{t}|X_{0})$ represents the process of gradually adding noise from $X_{0}$ to $X_{t}$, and $f_{i}(X_{0})$ is a $i$-th function representing the addition of noise to $X_{i-1}$ to obtain $X_{i}$. For example, in Figure 3, $f_{1}$ represents the operation of reverting the modification to the $O_{1}$ block in the evolved code, indicating that there is still a part of the code that needs to be modified. Each $X_{t}$ represents the state of the code at a specific stage of evolution, providing explicit direction that helps the model approach the final state at different modification stages (i.e., constructing the input-output pairs like $X_{t}$, $X_{0}>$). This strategy improves the model’s adaptability to diverse contexts and enhances its ability to capture subtle changes in the code.
@@
Existing diffusion models for discrete text generation typically add and remove noise at the continuous embedding level step by step. While this can increase the diversity of the generated text, it requires more iterative steps, making the process slower compared to the mainstream auto-regressive text generation approaches, resulting in lower efficiency. For example, in the upper part of Figure 2, each denoising step in traditional diffusion models modifies the overall embedding, allowing every token to potentially change in the next step, which facilitates generating diverse text. In contrast, in autoregressive generation, once a token is generated, its content and position are fixed, and the generation continues sequentially from that point onward, facilitating semantic coherence in the output [transformer2017, codet52021]. Therefore, we conduct the denoising process within an auto-regressive manner similar to CodeT5, using any intermediate state $X_{t}$ from the diffusion process as input and producing the final target $X_{0}$ as output. This approach maximizes the generative performance benefits of auto-regression. The denoising process can be formalized as:
@@
, where $p_{\theta}(X_{t-1}|X_{t})$ represents the probability distribution of denoising from $X_{t}$ to $X_{t-1}$, and $p_{\theta}(X_{t-1}^{i}|X_{t-1}^{1:i-1};X_{t})$ represents the probability of generating the $i$-th token in $X_{t-1}$, given the existing tokens in $X_{t}$ and generated tokens in $X_{t-1}$. The core characteristic of directional diffusion is reflected in each generation process of the auto-regressive model, where the generation target at any step is the final evolution goal. In other words, we transform intermediate diffusion data into input-output pairs for the auto-regressive model, such as $X_{T+1}$, $X_{0}>$, $X_{T}$, $X_{0}>$, $X_{T-1}$, $X_{0}>$. Under this strategy, each sample in the generation process is directional and optimized towards the final new code, thereby enhancing the model’s accuracy and consistency in generating outputs. Notably, by converting the directional diffusion process into input-output pairs for the pre-training task, our pre-trained model has the same inference cost as auto-regressive models like CodeT5 (requiring a single inference pass), which is lower compared to the higher inference cost of traditional diffusion models (requiring multiple inference passes).
@@ III-D Directional Diffusion Pretraining Tasks
DivoT5 is designed to acquire real-world code editing knowledge by learning to transform old code into completely new code through four directional diffusion-based pre-training tasks. To better illustrate the pre-training tasks, let us assume that the model only has the old code x ($x^{1},x^{2},\cdots,x^{M}$) as the input and the new code y ($y^{1},y^{2},\cdots,y^{N}$) as the output, where $M$ represents the number of input tokens and $N$ represents the number of output tokens. Figure 3 shows an illustration of the four tasks, where the corrupted code on the left ($X_{T+1}$) represents examples of the first three tasks, and the intermediate states $X_{t}$ ($0<t\leq T$) from the old code ($X_{T}$) in the middle to the new code ($X_{0}$) on the right illustrates the input for the last task. For clarity, we refer to the input code as the old code and the output code as the new code, and use blue italics to highlight the parts of the code that persist throughout the entire editing process. The first three tasks are designed to learn the basic editing processes and contextual dependencies of the code during the diffusion process, while the final task aims to reinforce the direction of code evolution. We place an additional special token at the prefix of the input sequence to distinguish between different tasks during the training process. Below we discuss the details of these tasks.
@@
This task aims to make the model aware of which code tokens do not need to be modified (i.e., Keep Span) during the code editing process. The design of this pre-training task leverages the observation that real-world code edits often retain a significant portion of the original structure and content. Since the code before and after evolution typically follows similar patterns or structures, some tokens remain unchanged across versions. For example, as shown in $X_{T+1(a)}$ of Figure 3, $[MASK1]$ represents the keyword that defines a variable (i.e., ‘public’), which commonly remains unchanged during edits. To leverage this property, we introduce a new masking strategy that explicitly masks unchanged tokens from real code edits, grounding the learning objective in the evolutionary direction. Unlike generic masking strategies over static code, our task is applied to real-world edit pairs, where each instance reflects a concrete transformation from an earlier version to a later one. By anchoring the keep span mask in this real edit context (i.e., evolutionary direction), we encourage the model to learn not only what to change but also what to preserve. This provides a valuable complementary learning signal: recognizing stable code regions is essential for generating accurate, minimal edits and avoiding unnecessary modifications.
Following existing works [cu2020, coditt52022], we calculate the differences between the new code and the old code. However, unlike existing techniques [lin2023cct5, codereviewer], our difference information is at the token level rather than at the line level. Then, we apply the SMLM [smlm2021] task specifically to the tokens marked as “KEEP” in the difference information. The formal representation of this task is as follows:
@@
, where $i$ represents the current generation step, $y_{i}$ represents the $i$-th token in the new code $y$. The $mask_{KSM}$ represents a type of SMLM approach, $x_{keep}$ indicates the unchanged parts of the code before and after editing, and $x_{change}$ indicates the modified parts. Since the number of unchanged tokens in this task is smaller than the total number of tokens in the old code, we mask 30% of the unchanged tokens, aiming to achieve a similar total number of masked tokens as in other tasks.
@@
This pre-training task treats the code editing process as a text generation task with evolutionary direction. Unlike the previous task, which focuses on unchanged regions, it treats all tokens equally by randomly masking token spans in the old code and learning to generate the completely new code through a sequence-to-sequence framework. Based on the characteristics of the dataset and existing works [codereviewer, coditt52022], we set the average span count to 2.5 and the overall percentage of masked tokens to be 20%. Unlike self-supervised tasks in the field of general text generation, the goal of this pre-training task is not simply to predict masked words, but rather to generate edited code in the presence of masks (i.e., prediction of the evolutionary direction). Formally, the loss of this task can be described as follows:
@@
Our third pre-training task combines the characteristics of denoising auto-encoding with the real-world code editing scenarios. It aims to train the model to generate the new code accurately while taking into account the noise and variations that occur in old code. Compared to the previous two pre-training tasks, this task demands a stronger ability for the model to understand code editing. However, this task well aligns with real-world development scenarios because we often cannot guarantee the quality of the old code. It reflects the practical nature of code editing, where the quality and effectiveness of the existing code vary. Specifically, we apply three types of noise to the old code: insertion, deletion, and replacement [coditt52022]. As shown in the code snippet $X_{T+1(c)}$ of Figure 3, in the old code, we add a $[MASK0]$ token, delete the keyword tokens like publicpublic, and replace specific variable names with $[MASK1]$. These modifications introduce noise and variations to simulate real-world code editing scenarios. During the generation process, the model not only needs to consider the tokens replaced by $[MASK]$, but also needs to take into account the scenario where existing tokens are deleted and the case of useless masks. Unlike existing techniques[coditt52022, li2023codeeditor] that utilize DAE with controlled noises to restore the old code, our approach requires the model to generate the actual new code after removing the introduced noise. The loss function for this pre-training task is as follows: